OpenStudy (anonymous):

Write y=logx+100 in the form y=Af(B(x-h))+k Identify A, B, h, K and describe the transformation

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@greenglasses

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large y = \log x +100\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\[\large y = 1\cdot \log( 1(x-0)) +100\]

ganeshie8 (ganeshie8):

compare that with the form \[\large y = A\cdot f( B(x-h)) +k\]

ganeshie8 (ganeshie8):

\(\large A = ?\) \(\large B = ?\) \(\large h = ?\) \(\large k=?\)

OpenStudy (anonymous):

okay so one example of it would be y=1/2x^2+6x+9 --> y=1/2(x+6)^2-9 its like vertex form

OpenStudy (anonymous):

so A=1, B=1, h=0, k=100

ganeshie8 (ganeshie8):

looks good!

ganeshie8 (ganeshie8):

h=0, so no horizontal shift

ganeshie8 (ganeshie8):

k=100, what does this tell you ?

OpenStudy (anonymous):

okay then can you help me with this one: switch this one too: y=e^-2x+10

ganeshie8 (ganeshie8):

please stick to one question per post

ganeshie8 (ganeshie8):

you can ask a new question by closing this and starting a new post :)